# Large AI Models for Wireless Physical Layer

### v1

@@ V Future Works and Open Issues
V-A Efficient and Lightweight Wireless LAMs
A significant barrier to the widespread adoption of LAMs in wireless systems is their substantial computational and memory requirement. Physical layer processing, particularly at the user side, is subject to stringent constraints on power consumption and latency. The massive parameter counts and complex computations inherent in current LAMs are often prohibitive for such environments. Therefore, a crucial line of future research is the development of efficient and lightweight wireless LAMs. This can be pursued through several avenues. First, model compression techniques, which have been successful in other domains, should be adapted for wireless LAMs. This includes methods like knowledge distillation, parameter pruning, and quantization. Second, exploring novel and efficient model architectures designed for LAMs is essential. Architectures like Mixture-of-Experts (MoE), which only activate a fraction of the model’s parameters for any given input, or State-Space Models (SSMs) that promise linear scaling complexity, could offer compelling alternatives to the NN blocks like the standard Transformer.
@@ V-B Interpretability of Wireless LAMs
For LAMs to be trusted and deployed in mission-critical communication networks, their decision-making processes cannot remain opaque ”black boxes.” Understanding why a model makes a particular prediction is fundamental for debugging, ensuring robustness, and certifying reliability. Consequently, developing eXplainable AI (XAI) techniques specifically for wireless LAMs is a paramount future task. Future work should focus on adapting and creating interpretability methods that are meaningful in the context of wireless communications. For instance, visualizing the attention mechanisms within Transformer-based LAMs can reveal which parts of the CSI data the model deems most important for a given task. This could provide valuable physical insights into the model’s learned strategies. Furthermore, techniques that assess feature importance, like LIME or SHAP, could be used to quantify the influence of different channel characteristics on the model’s output. A more profound goal would be to probe the internal representations of the LAM to see if they correspond to known physical phenomena (e.g., multipath clusters, angle of arrival/departure). Success in this area would not only build confidence in the models but could also lead to new, AI-driven discoveries about the nature of wireless channels.
@@ V-C Standardized Large-Scale Wireless Datasets
The progress in LAMs for NLP and CV was profoundly accelerated by the availability of massive, standardized benchmark datasets like Common Crawl and ImageNet. The wireless community currently lacks an equivalent resource, which hinders reproducible research and fair comparison of different models. The creation of a truly large-scale, comprehensive, and standardized wireless dataset is arguably one of the most important infrastructural tasks for the future of this field. This future dataset must go beyond the scale of existing ones. It should be characterized by:
  • Massive Scale and Diversity: Containing billions of channel samples covering an extensive range of scenarios (e.g., dense urban, indoor, rural, high-speed mobility), carrier frequencies (from Sub-6GHz to mmWave and THz), and antenna configurations (from massive MIMO to extremely large-scale arrays).
  • Hybrid Data Sources: Integrating high-fidelity data from both state-of-the-art simulation platforms (e.g., ray-tracing) and large-scale real-world measurement campaigns.
  • Rich Metadata: Crucially, the dataset should not only contain raw channel data but also be augmented with rich metadata, including precise location information, environmental maps (e.g., 3D building models), and corresponding data from other sensors (e.g., LiDAR, camera images) where available.
Such a resource would not only serve as a standard benchmark for training and evaluating wireless LAMs but also catalyze research into multimodal models that can achieve a more holistic understanding of the communication environment.
V-D Real-Time Inference and Hardware Acceleration
Beyond model size, the sheer computational complexity of LAMs poses a direct challenge to their deployment in real-world networks, where physical layer operations must often be completed within sub-millisecond timeframes. Bridging the gap between the complex processing required by LAMs and the strict latency demands of wireless communications is a critical challenge that requires synergistic innovation in algorithms and hardware. Future research must focus on optimizing LAMs for low-latency inference. This includes developing specialized inference strategies and redesigning model components, such as the self-attention mechanism, to reduce their computational complexity. However, algorithmic optimizations alone are unlikely to suffice. A pivotal direction will be the algorithm-hardware co-design of LAMs and the platforms they run on. This involves creating specialized hardware accelerators, such as application-specific integrated circuits (ASICs) and field-programmable gate arrays (FPGAs), that are purpose-built to efficiently execute the core operations of wireless LAMs. By designing the model architecture with hardware capabilities in mind and tailoring the hardware to the model’s computational patterns, it will be possible to dramatically reduce inference times. Achieving real-time performance is not merely an optimization problem but a fundamental prerequisite for translating the theoretical potential of wireless LAMs into practical reality.
V-E Large and Small AI Model Collaboration
@@ VI Conclusion
This paper reviews the application of LAMs to the wireless physical layer, highlighting their robust generalization, multitask processing, effective few-shot learning, and advanced multimodal capabilities. These attributes enable LAM-based approaches to achieve competitive performance with superior generalization in future communication systems. Despite this potential, the integration of LAMs into the physical layer remains an emerging field, requiring further exploration. Comprehensive theoretical analyses, practical deployment strategies, and innovative LAM architectures are essential to realize this paradigm in real-world communication scenarios.

### v2

@@ IV-C2 Universal LAM for Wireless Physical Layer
IV-D Quantitative Complexity Analysis
The complexity of wireless LAMs is a primary deployment concern. To provide a concrete view, we analyze the CSI feedback LAM [5] as an exemplar. Table II summarizes the complexity across Transformer-based models that differ in the number of blocks. We quantify complexity through per-inference CPU/GPU latency, NN parameter numbers, and floating-point operations (FLOPs).
From Table II, increasing the number of Transformer blocks in LAMs substantially raises both parameters and FLOPs. For example, relative to the tiny model, the LAM’s FLOPs are approximately $11.46\times$ higher, and the CPU runtime increases from $3.36$ ms to $20.53$ ms. This trend does not persist on the GPU: leveraging parallelism, the LAM’s runtime rises only from $0.90$ ms to $2.55$ ms, underscoring the efficiency of GPU inference for large-scale NNs. In this case, the modest GPU-side latency increase accompanies a $43\%$ reduction in feedback error compared with the tiny model, which represents a pragmatic trade-off for deployment.
@@ V Future Works and Open Issues
V-A Efficient, Real-Time, and Hardware-Aware Wireless LAMs
Both pretrained LAMs and wireless-native LAMs face tight latency, power, and memory budgets at the physical layer, especially on the user side where tasks such as CSI processing, beam prediction, and detection often require submillisecond inference under limited energy. Future work should focus on three concrete fronts: compress and right‑size models; engineer low‑latency inference with efficient dataflow; and align algorithms with accelerators under actual device limits. First, the model size should be reduced through compression and compact designs, using practical distillation, pruning, quantization, and sparsity to reduce parameters and FLOPs while preserving accuracy across diverse scenarios. Second, inference should be accelerated by simplifying core components, organizing dataflow to minimize memory movement, and applying lightweight attention, caching, or early exit where appropriate so execution aligns with physical layer timing. Third, algorithms should be co-designed with accelerators so that precision, operators, and scheduling match real device constraints on platforms like ASIC.
@@ V-B Interpretability of Wireless LAMs
For LAMs to be trusted and deployed in mission-critical communication networks, their decision-making processes cannot remain opaque ”black boxes.” Understanding why a model makes a particular prediction is fundamental for debugging, ensuring robustness, and certifying reliability. Consequently, developing eXplainable AI (XAI) techniques specifically for wireless LAMs is a paramount future task. Future work should focus on adapting and creating interpretability methods that are meaningful in the context of wireless communications. For instance, visualizing the attention mechanisms within Transformer-based LAMs can reveal which parts of the CSI data the model deems most important for a given task. This could provide valuable physical insights into the model’s learned strategies. Furthermore, techniques that assess feature importance could be used to quantify the influence of different channel characteristics on the model’s output. A more profound goal would be to probe the internal representations of the LAM to see if they correspond to known physical phenomena (e.g., multipath clusters, angle of arrival/departure). Success in this area would not only build confidence in the models but could also lead to new, AI-driven discoveries about the nature of wireless channels.
@@ V-C Standardized Large-Scale Wireless Datasets
The progress of LAMs in NLP and CV has been driven by massive, standardized datasets such as Common Crawl and ImageNet. The wireless community lacks an equivalent resource, hindering reproducibility and fair comparison. Building a truly large-scale, comprehensive, standardized wireless dataset is therefore a priority, and it must exceed the scale of existing efforts. This dataset should offer billions of channel samples across dense urban, indoor, rural, and high-speed scenarios, span Sub‑6GHz to upper 6 GHz and mmWave, and cover antenna configurations from massive MIMO to extremely large-scale arrays. It should combine high-fidelity outputs from state-of-the-art simulators (for example, ray tracing) with measurements from large real-world campaigns, and be enriched with metadata such as precise location, environmental maps (such as 3D building models), and, where available, LiDAR and camera data. Such a resource would provide a standard benchmark for training and evaluating wireless LAMs and catalyze multimodal research toward a holistic understanding of the communication environment.
V-D Large and Small AI Model Collaboration
@@ VI Conclusion
This paper reviews the application of LAMs to the wireless physical layer, highlighting their robust generalization, multitask processing, effective few-shot learning, and advanced multimodal capabilities that overcome limitations in conventional AI methods. These attributes enable LAM-based approaches to achieve competitive performance with superior generalization in future communication systems. Despite this potential, the integration of LAMs into the physical layer remains an emerging field, requiring further exploration. Comprehensive theoretical analyses, practical deployment strategies, and innovative LAM architectures are essential to realize this paradigm in real-world communication scenarios.